Please help me solve this equation. I don't know what to do.
(-1/2)(-2/3)
Thank you,
Cindy

Answer by stanbon(48535)

(-1/2)(-2/3)
That is not an equation.
The product, however, is [-1*-2]/[2*3] = 2/6 = 1/3
=================
Cheers,
Stan H.